# Test-Time Alignment of LLMs via Sampling-Based Optimal Control in pre-logit space

### v1

@@ 3.4 Modeling pre-logits distributions by Gaussian distributions
How the Gaussian assumption simplifies the problem
As described in Theorem 3.2, the Gaussian assumption derives a simple optimization procedure by using importance sampling. This is because the KL divergence between the optimal distribution and the Gaussian distribution is minimized by the expectation of $\bm{v}_{t}$ over $\mathbb{Q}^{*}$, and the weight function becomes simple because input trajectories also follow the Gaussian distribution. If we impose no constraints on the prior distribution, this computation requires modeling techniques for complicated distributions such as normalizing flows (Power & Berenson, 2023) and does not yield a simple test-time alignment.
@@
The softmax function is derived when the conditional distribution of pre-logits $p(\bm{z}_{t}|y_{t}\!=\!y^{i})$ is an exponential family distribution. [^2] This section considers a conditional distribution given an output token $y_{t}$ not given the past tokens $\bm{y}_{<t}$. If $p(\bm{z}_{t}|y_{t}\!=\!y^{i})$ is a Gaussian distribution as $p(\bm{z}_{t}|y_{t}\!=\!y^{i})\!=\!\mathcal{N}(\bm{\mu}_{y^{i}},\bm{\Sigma})$, we have the following equality from Bayes’ theorem:
$\textstyle P(y_{t}=y^{i}|\bm{z}_{t})=$ $\textstyle\frac{p(\bm{z}_{t}|y_{t}=y^{i})P(y_{t}=y^{i})}{p(\bm{z}_{t})}\textstyle=\frac{p(\bm{z}_{t}|y_{t}=y^{i})P(y_{t}=y^{i})}{\sum_{j}^{|\mathcal{V}|}p(\bm{z}_{t}|y_{t}=y^{j})P(y_{t}=y^{j})}$ (16)
$=$ $\textstyle\frac{\mathrm{exp}(\bm{\mu}_{y^{i}}^{T}\Sigma^{-1}\bm{z}-\frac{1}{2}\bm{\mu}_{y^{i}}\Sigma^{-1}\bm{\mu}_{y^{i}}+\ln P(y_{t}=y^{i}))}{\sum_{j}^{|\mathcal{V}|}\mathrm{exp}(\bm{\mu}_{y^{j}}^{T}\Sigma^{-1}\bm{z}-\frac{1}{2}\bm{\mu}_{y^{j}}\Sigma^{-1}\bm{\mu}_{y^{j}}+\ln P(y_{t}=y^{j}))}.$ (17)
This function corresponds to the softmax function with a linear layer such that $\bm{w}_{i}\!=\!\bm{\mu}_{y^{i}}^{\top}\Sigma^{-1}$ and $\bm{b}_{i}\!=\!-\frac{1}{2}\bm{\mu}_{y^{i}}\Sigma^{-1}\bm{\mu}_{y^{i}}\!+\!\ln P(y_{t}=y^{i})$ . From this result, Lee et al. (2018) assume that the pre-trained neural classifier has the pre-logtis following the Gaussian distribution given a class label in image recognition. Since neural language models also use softmax and cross-entropy loss, we can hypothesize that the trained pre-logit distribution $p(\bm{z}_{t}|y_{t})$ follows a Gaussian distribution. From this perspective, AISP can be regarded as exploring the optimal pre-logit distribution $p(\hat{\bm{z}}_{t}|y_{t}^{*})$ under the assumption, where the distribution given the optimal response can be decomposed as $p(\hat{\bm{z}}_{1},\dots,\hat{\bm{z}}_{t}|\mathbf{y}^{*})=\prod_{t}p(\hat{\bm{z}}_{t}|y_{t}^{*})$.
@@ 3.5 Connection with BoN
From this result, AISP can be regarded as a continuous approximation of BoN with a specific sampling strategy. In other words, AISP is a generalization of BoN, and AISP subsumes BoN.
@@ 3.6 Implementation
As discussed above, $\lambda$ can be regarded as a temperature parameter. Small $\lambda$ allows deviation from the base LLM, but large $\lambda$ causes numerical instability (Williams et al., 2018). To achieve both numerical stability and large penalties, we relax $\mathbb{P}$ as $p(V|\alpha\hat{U},\sigma)$ from $p(V|\bm{0},\sigma)$ where $0\!<\!\alpha\!<\!1$ by introducing the technique in MPPI (Williams et al., 2018). Under this relaxation, the weight $\bar{w}^{i}$ becomes
$\bar{w}^{i}=$ $\frac{\mathrm{exp}\left(\frac{1}{\lambda}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V^{i}))-\frac{1-\alpha}{\sigma^{2}}\sum_{t=1}^{\tau}\hat{\bm{u}}_{t}^{\top}\bm{v}^{i}_{t}\right)}{\sum_{j}\mathrm{exp}\left(\frac{1}{\lambda}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V^{j}))-\frac{1-\alpha}{\sigma^{2}}\sum_{t=1}^{\tau}\hat{\bm{u}}_{t}^{\top}\bm{v}^{j}_{t}\right)}.$ (19)
@@
Figure 2: Schematic illustration of computational costs (vertical: parallelism; horizontal: iterations) of Batched AISP and BoN for $D$ prompts. When $\kappa D/b\!=\!D,nb\!=\!N\Leftrightarrow\kappa\!=\!b,n\!=\!N/b$, Batched AISP and BoN have almost the same sequential and parallel computational cost.
@@ 4 Related work
There are several test-time alignment methods that train other networks to evaluate outputs or states of the base LLMs (Kong et al., 2024; Mudgal et al., 2024; Han et al., 2024; Kim et al., 2023). Critic-Guide Decoding (Kim et al., 2023) trains critic-networks that predict state-values of the current partial tokens. Similarly, Controlled Decoding (Mudgal et al., 2024) trains a value function and enables the evaluation of block-wise generation. RE-Control (Kong et al., 2024) also trains a value function but for optimizing the pre-logit. Though these methods avoid the high computational costs of fine-tuning LLMs, they still require the training process of value functions, and some need to build datasets in advance (Kong et al., 2024; Mudgal et al., 2024; Han et al., 2024). Among such methods, we compare AISP with RE-Control because it is the most similar to AISP. Khanov et al. (2024) proposes ARGS, which adds the weighted reward to the logit for each token. ARGS can be used as a training-free test-time alignment given a reward model, and we also compare AISP with it.
@@
As a sampling-based method, BoN is a popular method. Its improvement methods and performance analysis are actively investigated (Snell et al., 2024; Lightman et al., 2023; Brown et al., 2024; Ichihara et al., ; Jinnai et al., 2024). Snell et al. (2024) investigated the computational cost in test-time alignments, including BoN. They showed the difference between the characteristics of beam-search-based reward optimization and BoN: BoN outperforms beam-search when using a high computational budget. They concluded that there is a compute-optimal scaling strategy, which acts to most effectively allocate test-time compute adaptively per prompt. AISP can be included in such a strategy. While it is revealed that BoN’s win-rate against a base LLM is bounded by $N/(N+1)$ (Beirami et al., 2024), there are few studies to improve the efficiency of reward optimization in terms of the number of used samples. As another sampling method, Zhu et al. (2025) proposed Soft Reasoning based on Bayesian optimization. Soft Reasoning adds the Gaussian perturbation to a token embedding, similar to AISP, and applies Bayesian optimization. However, Soft Reasoning only applies the perturbation to the initial token embedding, which might be due to the difficulty in Bayesian optimization. Since AISP converges to the optimal distribution, it optimizes a more complicated pre-logit sequence than the initial token embedding. Loula et al. (2025) used importance sampling to control generation of LLMs on the basis of the given potential function, and extended it to enable the evaluation of the partial sequence during generation. Though it is similar to our method, they generate tokens using task-specific potential functions rather than the reward model. Additionally, they use importance sampling on the token space rather than the pre-logit space. Because of using pre-logits distributions, our method can employ a Gaussian distribution, which is easy to handle.
@@ 5.1 Setup
We conduct experiments to evaluate the effectiveness of AISP on test-time alignment of LLMs for helpfulness and minimizing harmfulness. We use Anthropic’s HH-RLHF (Bai et al., 2022) and Stanford human preferences (SHP) datasets (Ethayarajh et al., 2022) following (Kong et al., 2024). These datasets are used to align LLMs for helpfulness and harmlessness. We use randomly selected 1000 entries of the test datasets due to limited computation resources, like (Jinnai et al., 2024). We use Llama-3-8B (AI@Meta, 2024), Vicuna-7B-v1.5 (Chiang et al., 2023) [^3] https://huggingface.co/lmsys/vicuna-7b-v1.5, and Gemma3-4B (Team, 2025) as the base LLMs, and reward models are UltraRM-13b (UltraRM) (Cui et al., 2023) and Eurus-RM-7b (Eurus) (Yuan et al., 2024).
@@
We compare AISP with BoN using top-$p$ (nucleus) sampling. Both $n$ and $\kappa$ of AISP are set to 32, and $N$ of BoN is set to 1024 ($\kappa n$). Additionally, we also compare AISP with ARGS-greedy (ARGS) Khanov et al. (2024) and RE-Control Kong et al. (2024), which are reward-based test-time alignment methods. ARGS adds the weighted reward to logits and selects the best token for each token generation, and RE-Control adds the control input to maximize the value function. We tune hyper-parameters for each method on partial training datasets, which is described in Appendix C.2. Sensitivity to hyper-parameter in AISP is shown in Appendix D.1
@@
Llama3_8B BoN (top-$p$) -2.38 (0.04) 0.693 (0.009) 0.623 (0.004) -5.074 (0.007) 0.742 (0.002) 0.605 (0.007)
& UltraRM RE-Control -9.28 (0.03) 0.836 (0.003) 0.559 (0.004) -5.531 (0.009) 0.743 (0.08) 0.573 (0.006)
ARGS -3.94 0.786 0.531 -9.58 0.338 0.596
AISP -1.39 (0.02) 0.773 (0.004) 0.626 (0.004) -5.02 (0.01) 0.724 (0.000) 0.578 (0.001)
Vicuna_7B BoN (top-$p$) -1.78 (0.02) 0.882 (0.002) 0.658 (0.000) -4.85 (0.01) 0.804 (0.004) 0.615 (0.001)
& UltraRM RE-Control -5.67 (0.04) 0.843 (0.001) 0.654 (0.001) -5.25 (0.02) 0.610 (0.006) 0.527 (0.01)
ARGS -11.97 0.774 0.066 -8.37 0.614 0.092
AISP -1.46 (0.02) 0.884 (0.002) 0.654 (0.001) -4.73 (0.02) 0.803 (0.003) 0.599 (0.000)
Gemma3_4B BoN (top-$p$) -3.43 (0.02) 0.879 (0.003) 0.646 (0.002) -5.26 (0.005) 0.809 (0.002) 0.539 (0.008)
& UltraRM RE-Control -9.97 (0.02) 0.862 (0.001) 0.556 (0.005) -5.78 (0.02) 0.824 (0.007) 0.615 (0.02)
ARGS -7.08 0.910 0.192 -7.54 0.917 0.189
AISP -2.39 (0.03) 0.819 (0.008) 0.675 (0.003) -5.24 (0.03) 0.758 (0.005) 0.555 (0.003)
Llama3_8B BoN (top-$p$) -6.42 (0.08) 0.758 (0.000) 0.644(0.003) -5.07 (0.04) 0.736 (0.004) 0.669 (0.000)
& Eurus RE-Control -9.62 (0.1) 0.793 (0.02) 0.540 (0.01) -5.62 (0.05) 0.727 (0.01) 0.572 (0.01)
ARGS -11.91 0.585 0.425 -7.76 0.290 0.597
AISP -6.17 (0.03) 0.750 (0.006) 0.659 (0.004) -5.11 (0.02) 0.715 (0.007) 0.662 (0.003)
Vicuna_7B BoN (top-$p$) -3.83 (0.02) 0.884 (0.001) 0.654 (0.001) -4.88 (0.03) 0.800 (0.002) 0.648 (0.003)
& Eurus RE-Control -5.24 (0.03) 0.8555 (0.002) 0.653 (0.001) -5.46 (0.1) 0.529 (0.01) 0.475 (0.03)
ARGS -12.67 0.843 0.226 -8.31 0.791 0.146
AISP -3.72 (0.02) 0.896 (0.000) 0.651 (0.002) -4.85 (0.01) 0.806 (0.002) 0.648 (0.001)
Gemma3_4B BoN (top-$p$) -6.45 (0.06) 0.856 (0.001) 0.639 (0.004) -5.34 (0.008) 0.732 (0.007) 0.665 (0.002)
& Eurus RE-Control -10.1 (0.1) 0.853 (0.01) 0.552 (0.006) -5.82 (0.06) 0.817 (0.004) 0.607 (0.003)
ARGS -14.1 0.949 0.195 -7.53 0.917 0.189
AISP -5.78 (0.03) 0.814 (0.002) 0.656 (0.003) -5.38 (0.03) 0.794 (0.004) 0.643 (0.001)

### v2

@@ 3.4 Modeling pre-logits distributions by Gaussian distributions
How Gaussian assumption simplifies the problem
Theorem 3.2 indicates that the Gaussian assumption derives a simple optimization procedure by using importance sampling. This is because the KL divergence between the optimal and the Gaussian distributions is minimized by the expectation of $\bm{v}_{t}$ over $\mathbb{Q}^{*}$, and the weight function becomes simple because input trajectories also follow the Gaussian distribution. If we impose no constraints on the prior distribution, this computation requires modeling techniques for complicated distributions such as normalizing flows (Power and Berenson, 2023) and does not yield a simple method.
@@
The softmax function is derived when the conditional distribution of pre-logits $p(\bm{z}_{t}|y_{t}\!=\!y^{i})$ is an exponential family distribution. [^2] This section considers a conditional distribution given an output token $y_{t}$ not given the past token sequence $\bm{y}_{<t}$. If $p(\bm{z}_{t}|y_{t}\!=\!y^{i})$ is a Gaussian distribution as $p(\bm{z}_{t}|y_{t}\!=\!y^{i})\!=\!\mathcal{N}(\bm{\mu}_{i},\bm{\Sigma})$, we have the following equality from Bayes’ theorem:
$\textstyle P(y_{t}\!=\!y^{i}|\bm{z}_{t})\!=$ $\textstyle\frac{p(\bm{z}_{t}|y_{t}=y^{i})P(y_{t}=y^{i})}{p(\bm{z}_{t})}\textstyle$
$=$ $\textstyle\frac{p(\bm{z}_{t}|y_{t}=y^{i})P(y_{t}=y^{i})}{\sum_{j}^{|\mathcal{V}|}p(\bm{z}_{t}|y_{t}=y^{j})P(y_{t}=y^{j})}$
$=$ $\textstyle\frac{\mathrm{exp}(\bm{\mu}_{i}^{\top}\!\Sigma^{-1}\bm{z}-\frac{1}{2}\bm{\mu}_{i}^{\top}\!\Sigma^{-1}\bm{\mu}_{i}+\ln\!P(y_{t}=y^{i}))}{\sum_{j}^{|\mathcal{V}|}\mathrm{exp}(\bm{\mu}_{j}^{\top}\!\Sigma^{-1}\bm{z}-\frac{1}{2}\bm{\mu}_{j}^{\top}\!\Sigma^{-1}\!\bm{\mu}_{j}+\ln\!P(y_{t}=y^{j}))}\!.$ (16)
This function corresponds to the softmax function with a linear layer such that $\bm{w}_{i}\!=\!\bm{\mu}_{y^{i}}^{\top}\Sigma^{-1}$ and $\bm{b}_{i}\!=\!-\frac{1}{2}\bm{\mu}_{y^{i}}^{\top}\Sigma^{-1}\bm{\mu}_{y^{i}}\!+\!\ln P(y_{t}\!=\!y^{i})$. From this result, Lee et al. (2018) assume that the pre-trained neural classifier has the pre-logtis following the Gaussian distribution given a class label in image recognition. Since neural language models also use softmax and cross-entropy loss, we can hypothesize that the trained pre-logit distribution $p(\bm{z}_{t}|y_{t})$ follows a Gaussian distribution. From this perspective, AISP can be regarded as exploring the optimal pre-logit distribution $p(\hat{\bm{z}}_{t}|y_{t}^{*})$ under the assumption, where the distribution given the optimal response can be decomposed as $p(\hat{\bm{z}}_{1},\dots,\hat{\bm{z}}_{t}|\mathbf{y}^{*})\!=\!\prod_{t}p(\hat{\bm{z}}_{t}|y_{t}^{*})$. We evaluate the effectiveness of Gaussian assumption by using BoN with Gaussian perturbations, as explained in next section. Note that Gaussian assumption is commonly used in neural language model: Li et al. (2021) modeled embeddings by using Gaussian distributions, and Podolskiy et al. (2021) used Mahalanobis distance between embeddings, which can be derived from the Gaussian assumption.
@@ 3.5 Connection with BoN
$\textstyle\mathcal{Y}_{n}=\{\bm{\mathrm{y}}(V^{i})|V^{i}\sim q(V|\hat{U},\sigma^{2}),i=1,\dots,n\}.$ (17)
From this result, AISP can be regarded as a continuous approximation of BoN with a specific candidate set. In other words, AISP is a generalization of BoN, and AISP subsumes BoN. Comparing BoN using Eq. (17) with BoN using popular sampling strategies, we can evaluate the Gaussian assumtpion in AISP. Thus, we empirically evaluate BoN using this candidate set when $\hat{U}\!=\!\bm{O}$ and observed it achieves compatitive performance in experiments.
@@ 3.6 Implementation
As discussed above, $\lambda$ can be regarded as a temperature parameter. Small $\lambda$ allows deviation from the base LLM, but large $\lambda$ causes numerical instability (Williams et al., 2018). To achieve both numerical stability and large penalties, we relax $\mathbb{P}$ as $p(V|\alpha\hat{U},\sigma)$ from $p(V|\bm{0},\sigma)$ where $0\!<\!\alpha\!<\!1$ by using the technique in MPPI (Williams et al., 2018). As a result, the weight $\bar{w}^{i}$ becomes
$\textstyle\bar{w}^{i}=\frac{\mathrm{exp}\left(\frac{1}{\lambda}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V^{i}))-\frac{1-\alpha}{\sigma^{2}}\sum_{t=1}^{\tau}\hat{\bm{u}}_{t}^{\top}\bm{v}^{i}_{t}\right)}{\sum_{j}\mathrm{exp}\left(\frac{1}{\lambda}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V^{j}))-\frac{1-\alpha}{\sigma^{2}}\sum_{t=1}^{\tau}\hat{\bm{u}}_{t}^{\top}\bm{v}^{j}_{t}\right)}.$ (18)
@@
Figure 2: Schematic illustration of computational costs (vertical: parallelism; horizontal: iterations) of Batched AISP and BoN. When $\kappa D/b\!=\!D,nb\!=\!N\Leftrightarrow\!\kappa\!=\!b,n\!=\!N/b$, they have almost the same sequential and parallel computational cost.
@@ 4 Related work
There are several test-time alignment methods that train other networks to evaluate outputs or states of the base LLMs (Kong et al., 2024; Mudgal et al., 2024; Han et al., 2024; Kim et al., 2023). Critic-Guide Decoding (Kim et al., 2023) trains critic-networks that predict state-values of the current partial tokens. Similarly, Controlled Decoding (Mudgal et al., 2024) trains a value function and enables the evaluation of block-wise generation. RE-Control (Kong et al., 2024) also trains a value function but for optimizing the pre-logit. Xu et al. (2025) have presented an autoregressive reward model (GenARM) for the test-alignment. Though these methods avoid the high computational costs of fine-tuning LLMs, they still require the training process of value functions or autoregressive rewards, and some need to build datasets in advance (Kong et al., 2024; Mudgal et al., 2024; Han et al., 2024). Among such methods, we compare AISP with RE-Control because it is the most similar to AISP. (Khanov et al., 2024) proposes ARGS, which adds the weighted reward to the logit for each token. ARGS can be used as a training-free test-time alignment given a reward model, and we also compare AISP with it. Chakraborty et al. (2024) have presented transfer-$Q^{*}$ that estimates token-level value function through the trajectory-based reward function. When the base LLM is not aligned with the given target reward in advance, transfer-$Q^{*}$ requires the base reward model, and it is hard to fairly compare AISP with it.
@@
BoN is one of the most popular methods as sampling-based alignment. Its improvement methods and performance analysis are actively investigated (Snell et al., 2024; Lightman et al., 2023; Brown et al., 2024; Ichihara et al., ; Jinnai et al., 2024). Snell et al. (2024) investigated the computational cost in test-time alignments, including BoN. They showed the difference between the characteristics of beam-search-based method and BoN: BoN outperforms beam-search when using a high computational budget. They concluded that there is a compute-optimal scaling strategy, which acts to most effectively allocate test-time compute adaptively per prompt. AISP can be included in such a strategy. While it is revealed that BoN’s win-rate against a base LLM is bounded by $N/(N+1)$ (Beirami et al., 2024), there are few studies to improve the efficiency of reward optimization in terms of the number of used samples. As another sampling method, Zhu et al. (2025) proposed Soft Reasoning based on Bayesian optimization. Soft Reasoning adds the Gaussian perturbation to a token embedding, similar to AISP, and applies Bayesian optimization. However, Soft Reasoning only applies the perturbation to the initial token embedding, which might be due to the difficulty in Bayesian optimization. Since AISP converges to the optimal distribution, it optimizes a more complicated pre-logit sequence than the initial token embedding. Loula et al. (2025) used importance sampling to control generation of LLMs on the basis of the given potential function. Though it is similar to our method, they generate tokens using task-specific potential functions rather than the reward model. Additionally, they use importance sampling on the token space rather than the pre-logit space. Thanks to continuous pre-logits space, our method can employ a Gaussian distribution, which is easy to handle.
@@ 5.1 Setup
We conduct experiments to evaluate the effectiveness of AISP on test-time alignment of LLMs for helpfulness and minimizing harmfulness. We use Anthropic’s HH-RLHF (Bai et al., 2022) and Stanford human preferences (SHP) datasets (Ethayarajh et al., 2022) following (Kong et al., 2024). These datasets are used to align LLMs for helpfulness and harmlessness. We use randomly selected 1000 entries of the test datasets due to limited computation resources, like (Jinnai et al., 2024). We use Llama-3-8B (AI@Meta, 2024), Vicuna-7B-v1.5 (Chiang et al., 2023) [^3] https://huggingface.co/lmsys/vicuna-7b-v1.5, and Gemma3-4B (Team, 2025) as the base LLMs, and reward models are UltraRM-13b (UltraRM) (Cui et al., 2023) and Eurus-RM-7b (Eurus) (Yuan et al., 2024). In Appendix D.4, we evaluate AISP on Alpaca-Eval (Li et al., 2023), GSM8K (Cobbe et al., 2021), HumanEval (Chen et al., 2021), and TruthfulQA (Lin et al., 2022).
@@
We compare AISP with BoN using two generating strategies to construct candidate sets $\mathcal{Y}_{N}$: BoN (top-p) and BoN ($\mathcal{N}$). BoN (top-p) uses top-p (nucleus) sampling, and BoN ($\mathcal{N}$) injects $n\kappa$ Gaussian noises following $\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ into pre-logits as Eq. (17). After injection, we use greedy search for each injected pre-logits to construct $\mathcal{Y}_{N}$ of BoN ($\mathcal{N}$). Both $n$ and $\kappa$ of AISP are set to 32, and $N$ of BoN is set to 1024 ($\kappa n$). Additionally, we also compare AISP with ARGS-greedy (ARGS) (Khanov et al., 2024) and RE-Control (Kong et al., 2024), which are reward-based test-time alignment methods. ARGS adds the weighted reward to logits and selects the best token for each token generation, and RE-Control adds the control input to maximize the value function. We tune hyper-parameters for each method on partial training datasets, which is described in Appendix C.2. Sensitivity to hyper-parameter in AISP is shown in Appendix D.1
@@
Llama3_8B BoN (top-$p$) -2.38 (0.04) 0.693 (0.009) 0.623 (0.004) -2.62 (0.003) 0.743 (0.005) 0.637 (0.004)
& UltraRM BoN ($\mathcal{N}$) -2.30 (0.03) 0.752 (0.005) 0.618 (0.003) -2.60 (0.01) 0.763 (0.000) 0.633 (0.002)
RE-Control -9.28 (0.03) 0.836 (0.003) 0.559 (0.004) -7.54 (0.06) 0.781 (0.000) 0.619 (0.000)
ARGS -3.94 0.786 0.531 -9.27 0.747 0.605
AISP -1.39 (0.02) 0.773 (0.004) 0.626 (0.004) -2.45 (0.00) 0.797 (0.003) 0.623 (0.000)
Vicuna_7B BoN (top-$p$) -1.78 (0.02) 0.882 (0.002) 0.658 (0.000) -3.08 (0.00) 0.861 (0.002) 0.619 (0.001)
& UltraRM BoN ($\mathcal{N}$) -1.94 (0.01) 0.876 (0.000) 0.659 (0.001) -3.02 (0.01) 0.860 (0.002) 0.617 (0.000)
RE-Control -5.67 (0.04) 0.843 (0.001) 0.654 (0.001) -4.43 (0.05) 0.798 (0.003) 0.611 (0.001)
ARGS -11.97 0.774 0.066 -11.53 0.739 0.146
AISP -1.46 (0.02) 0.884 (0.002) 0.654 (0.001) -2.86 (0.02) 0.871 (0.002) 0.615 (0.003)
Gemma3_4B BoN (top-$p$) -3.43 (0.02) 0.879 (0.003) 0.646 (0.002) -2.60 (0.02) 0.719 (0.002) 0.648 (0.000)
& UltraRM BoN ($\mathcal{N}$) -3.29 (0.02) 0.772 (0.004) 0.684 (0.002) -2.35 (0.01) 0.776 (0.005) 0.637 (0.001)
RE-Control -9.97 (0.02) 0.862 (0.001) 0.556 (0.005) -9.00 (0.2) 0.347 (0.1) 0.393 (0.04)
ARGS -7.08 0.910 0.192 -11.11 0.919 0.260
AISP -2.39 (0.03) 0.819 (0.008) 0.675 (0.003) -2.18 (0.02) 0.772 (0.007) 0.632 (0.001)
Llama3_8B BoN (top-$p$) -6.42 (0.08) 0.758 (0.000) 0.644(0.003) -5.02 (0.03) 0.702 (0.002) 0.646 (0.002)
& Eurus BoN ($\mathcal{N}$) -7.22 (0.04) 0.759 (0.003) 0.645 (0.003) -4.91 (0.03) 0.722 (0.006) 0.645 (0.000)
RE-Control -9.62 (0.1) 0.793 (0.02) 0.540 (0.01) -11.28 (0.3) 0.282 (0.06) 0.084 (0.01)
ARGS -11.91 0.585 0.425 -10.50 0.526 0.538
AISP -6.17 (0.03) 0.750 (0.006) 0.659 (0.004) -4.96 (0.05) 0.759 (0.006) 0.639 (0.001)
Vicuna_7B BoN (top-$p$) -3.83 (0.02) 0.884 (0.001) 0.654 (0.001) -3.71 (0.00) 0.874 (0.000) 0.629 (0.003)
& Eurus BoN ($\mathcal{N}$) -4.09 (0.02) 0.872 (0.001) 0.656 (0.001) -3.77 (0.02) 0.853 (0.002) 0.631 (0.000)
RE-Control -5.24 (0.03) 0.8555 (0.002) 0.653 (0.001) -10.24 (0.2) 0.072 (0.1) 0.031 (0.000)
ARGS -12.67 0.843 0.226 -11.46 0.850 0.234
AISP -3.72 (0.02) 0.896 (0.000) 0.651 (0.002) -3.65 (0.04) 0.872 (0.000) 0.620 (0.001)
Gemma3_4B BoN (top-$p$) -6.45 (0.06) 0.856 (0.001) 0.639 (0.004) -4.82 (0.05) 0.710 (0.005) 0.658 (0.001)
& Eurus BoN ($\mathcal{N}$) -6.30 (0.04) 0.774 (0.002) 0.644 (0.004) -4.84 (0.03) 0.747 (0.008) 0.654 (0.002)
[…]
ARGS -14.1 0.949 0.195 -11.74 0.933 0.025
AISP -5.78 (0.03) 0.814 (0.002) 0.656 (0.003) -4.80 (0.02) 0.800 (0.003) 0.648 (0.003)